OpenStudy (anonymous):

3x^2+2x-8 factor please help !!!!!!!!!!!!!

OpenStudy (campbell_st):

well here is a method that works for a quadratic \[ax^2 + bx + c = 0\] when you compare it to your equation, a = 3, b = 2 and c = -8 multiply a and c so 3 x -8 = -24 find the factors of -24 that add to b, which is 2.... the larger factor is positive. any thoughts..?

OpenStudy (anonymous):

Looking at 3x%5E2-2x-8 we can see that the first term is 3x%5E2 and the last term is -8 where the coefficients are 3 and -8 respectively. Now multiply the first coefficient 3 and the last coefficient -8 to get -24. Now what two numbers multiply to -24 and add to the middle coefficient -2? Let's list all of the factors of -24: Factors of -24: 1,2,3,4,6,8,12,24 -1,-2,-3,-4,-6,-8,-12,-24 ...List the negative factors as well. This will allow us to find all possible combinations These factors pair up and multiply to -24 (1)*(-24) (2)*(-12) (3)*(-8) (4)*(-6) (-1)*(24) (-2)*(12) (-3)*(8) (-4)*(6) note: remember, the product of a negative and a positive number is a negative number Now which of these pairs add to -2? Lets make a table of all of the pairs of factors we multiplied and see which two numbers add to -2 First Number Second Number Sum 1 -24 1+(-24)=-23 2 -12 2+(-12)=-10 3 -8 3+(-8)=-5 4 -6 4+(-6)=-2 -1 24 -1+24=23 -2 12 -2+12=10 -3 8 -3+8=5 -4 6 -4+6=2 From this list we can see that 4 and -6 add up to -2 and multiply to -24 Now looking at the expression 3x%5E2-2x-8, replace -2x with 4x%2B-6x (notice 4x%2B-6x adds up to -2x. So it is equivalent to -2x) 3x%5E2%2Bhighlight%284x%2B-6x%29%2B-8 Now let's factor 3x%5E2%2B4x-6x-8 by grouping: %283x%5E2%2B4x%29%2B%28-6x-8%29 Group like terms x%283x%2B4%29-2%283x%2B4%29 Factor out the GCF of x out of the first group. Factor out the GCF of -2 out of the second group %28x-2%29%283x%2B4%29 Since we have a common term of 3x%2B4, we can combine like terms So 3x%5E2%2B4x-6x-8 factors to %28x-2%29%283x%2B4%29 So this also means that 3x%5E2-2x-8 factors to %28x-2%29%283x%2B4%29 (since 3x%5E2-2x-8 is equivalent to 3x%5E2%2B4x-6x-8) ------------------------------- Answer: So 3x%5E2-2x-8 factors to %28x-2%29%283x%2B4%29

OpenStudy (campbell_st):

wow... so using what @AllisonRoyal posted the factors are 6 and - 4 so write the factorisation as \[\frac{(ax + factor 1)(ax + factor 2)}{a}~~~or~~~\frac{(3x +6)(3x - 4)}{3}\] now removed any common factors in the numerator and you'll see they(it) it will cancel with the denominator... a fool-proof method

OpenStudy (campbell_st):

and it avoids a lot og guess and check work

OpenStudy (anonymous):

thanks guys